Assume a point charge Q1 is at x=0 and a charge Q2 is at x= d. What is the relationship between these charges if the resulting electric field is zero at the following points:

a) x= d/2; b) x= 2d; c) x= -d/2


1
Expert's answer
2022-02-18T08:22:35-0500

First let us remind the expression of the total field at the point "x" :

"E= \frac{1}{4\pi\epsilon_0} (\frac{Q_1 x}{|x|^3}+\frac{Q_2 (x-d)}{|x-d|^3})"

  1. The first case gives us "\frac{Q_1}{d^2/4}-\frac{Q_2}{d^2/4}=0" which yields "Q_1=Q_2"
  2. The second case gives "\frac{Q_1}{4d^2}+\frac{Q_2}{d^2}=0" which yields "Q_1=-4Q_2"
  3. The third case gives "\frac{-Q_1}{d^2/4}+\frac{-Q_2}{9d^2/4}=0" which yields "Q_1=-Q_2\/9"